A man is riding a vehicle towards a vertical cliff with a velocity of 36 km/h. He fires a gunshot and hears its echo after 4 sec. The speed of sound in air is 320 m/s.

(i) The distance between the cliff and the position, where he shot the gun is:

A.640 m

B.680 m

C.620 m

D.660 m


1
Expert's answer
2021-10-18T10:59:32-0400

36 km/h is 10 m/s.

The speed of sound is added to the speed of the motorcycle when the sound moves toward the cliff, reflects from the cliff and travels back:


"t_1=d/(c+v),t_2=d/(c+v),\\
t=t_1+t_2=4\text{ s}.\\
d=t(c+v)/2=660\text{ m}."